An alpha particle with initial spee d v 0 is projected from infinity and it approaches up to r 0 distance from the nuclus. Then the initial speed of alpha particle which approaches upto 2r 0 distance from the nuclus. Is

An  alpha particle with initial speed v0 is projected from infinity and it approaches up to
r0 distance from the nuclus. Then the initial speed of alpha particle which approaches upto 2rdistance from the nuclus. Is

Grade:11

1 Answers

Khimraj
5 years ago
Kinetic energy  = potential energy
½*m*vo2 = k*2e*Ze/ro           – (1)
and let velocity is v
then
½*m*v2 = k*2e*Ze/2ro           -(2)
then (2)/(1)
v2/vo2 = ½
then v = vo/\sqrt{2}
Hope it clears.